Circular No. 8840

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 2007 K5 T. Lovejoy, Thornlands, Queensland, Australia, reports his discovery

of a comet on twelve 90-s CCD images obtained on May 26 with a 200-mm f/2.8 camera lens, the object reported as circular, 1′ across, with a clear blue-green color but no tail. J. Drummond (Gisborne, New Zealand, 0.20m f/4.6 reflector) writes that his confirming images from May 28, obtained at altitude ∼ 15^{o}, show a moderately condensed coma of diameter ∼ 1′ and no tail. The following reported positions include precise astrometry by R. H. McNaught from images obtained with the 0.5-m Uppsala Schmidt telescope at Siding Spring, who reported a circular coma 0′.6 in diameter. Additional accurate astrometry is needed.

2007 UT  α_{2000}  δ_{2000}  Mag. Observer
May 26.3424 5^{h}50^{m}24.^{s}9 −15^{o}09′16′′  13  Lovejoy
26.3458 5 50 25.4 −15 09 11  _{”}
28.28  6 00 18.2 −15 05 39.7 13  Drummond
28.31  6 00 24.87 −15 05 42.6  _{”}
28.3458 6 00 34.4 −15 04 56  13  Lovejoy
28.3556 6 00 34.5 −15 05 07  _{”}
29.27 6 05 15.77 −15 03 34.6 Drummond
29.34201 6 05 38.82 −15 03 15.2  McNaught
29.34442 6 05 39.56 −15 03 15.0  _{”}
29.34673 6 05 40.25 −15 03 14.8 13.8  _{”}

(88611) TEHARONHIAWAKO I (SAWISKERA) The companion to minor planet (88611) Teharonhiawako = 2001 QT_{297}

(cf. IAUCs 7733 and 7765) has been named Sawiskera (cf. MPC 59388).

COMET C/2006 P1 (McNAUGHT) Total visual magnitude estimates by S. Rae, Hamilton, New Zealand:

Mar. 9.38 UT, 6.2 (10×50 binoculars); 18.65, 6.5; 20.64, 6.7; 23.64, 8.4 (0.25-m reflector); 26.65, 8.7; Apr. 9.31, 9.3; 15.31, 9.5; 18.68, 9.6; 22.36, 9.7; 24.47, 9.6; May 12.35, 10.1.

COMET 96P/MACHHOLZ Total visual magnitudes by J. Gonzalez, Leon, Spain: Apr. 13.18 UT,

7.2 (25×100 binoc.); 18.16, 8.2; 22.17, 8.8; May 9.05, 10.9 (0.20-m refl.).

2007 May 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green